OpenStudy (kj4uts):

23° 20' 48" is the same as ____ °. Round to the nearest hundredth of a degree. Please explain.

OpenStudy (kj4uts):

Is my math correct here (Is this how you are supposed to solve these problems)?: 23(degrees)+(20/60)+(48/3600)=23.34666667 rounded to nearest hundredth of a degree would be: ANSWER: 23.35

OpenStudy (paxpolaris):

there are 60 minutes in a degree ... and there are 60 seconds in a minute...

OpenStudy (paxpolaris):

correct !

OpenStudy (kj4uts):

So the answer would be 23.35?

OpenStudy (paxpolaris):

yes

OpenStudy (kj4uts):

Thanks for your help and time!

OpenStudy (paxpolaris):

np

OpenStudy (kj4uts):

How would I solve this one though?

OpenStudy (kj4uts):

42.54° is the same as 42° 32' ____".

OpenStudy (freckles):

you could do this: we have 42 deg 32 min x sec \[(42+\frac{32}{60}+\frac{x}{60^2})^o=42.54^o\] solve for x

OpenStudy (kj4uts):

then the answer would be 24??

OpenStudy (freckles):

sounds good